OpenStudy (anonymous):

HOW MANY NUMBER CONSISTING OF TWO DIGITS CAN BE FORMED FROM 2,3,5,7.EACH INTEGER IS TO BE USED ONLY ONCE?

OpenStudy (anonymous):

12

OpenStudy (angela210793):

4*3=12

OpenStudy (anonymous):

HOW?

OpenStudy (angela210793):

for the 1st digit you have 4no.s u can put and for the 2nd one u have only 3 cause u can't put the one u have put in the 1st 1 :) Get it?

OpenStudy (anonymous):

SORRY.CAN U PLZ DO IT IN STEPS?

OpenStudy (angela210793):

w8 :)

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

ANGELA?????????????????

OpenStudy (angela210793):

a sec plsssss

OpenStudy (angela210793):

for the 1-st digit u can choose one of the no.s given(2;3;5;7) and for the 2nd digit are left only three no.s from which u can choose to put in there cause you can't put the same didgit twice EX: for the 1st digit you choose to put 3 and then for the 2nd digit you have left only 2,5,7 so 4*3=12

OpenStudy (amistre64):

4 Pick 2; 4P2 = 4*3 = 12

OpenStudy (amistre64):

4!/2! = 4P2 = 12 as well :)

OpenStudy (angela210793):

See the attcch. maybe u'll understand